How to express a 3rd order differential equation into three first order differential equations?

Explain.


1
Expert's answer
2021-06-06T12:59:24-0400

Let's 3rd order differential equation is "y'''(t) = f(t, y, y', y'')". let introduce new vector-function "\overrightarrow{x}(t)"

"\overrightarrow{x} = \begin{bmatrix}
 x_1(t) \\
 x_2(t) \\
 x_3(t)
\end{bmatrix} = \begin{bmatrix}
 y(t) \\
 y'(t) \\
 y''(t)
\end{bmatrix}"

then

"\frac{d \overrightarrow{x}}{d t} = \frac{d }{d t} 
\begin{bmatrix}
 y(t) \\
 y'(t) \\
 y''(t)
\end{bmatrix} =
\begin{bmatrix}
 x_2(t) \\
 x_3(t) \\
 f(t, x_1(t), x_2(t), x_3(t)
\end{bmatrix}"